# Quantum-Informed Machine Learning for Chaotic Systems

### v2

@@ IV.1 Quantum-informed machine learning framework for chaotic dynamical systems
Here, is the empirical distribution induced by predicted states, denotes a kernel embedding, and are empirically selected hyperparameters. The KL divergence captures low-order, information-theoretic discrepancies, while the MMD term captures higher-order statistical structure through kernel embeddings. The full QIML framework is demonstrated in Algorithm 1. All models are implemented in PyTorch with fully differentiable integration of Q-priors. Architectural details for predicting 1D and 2D systems are provided in Supplementary Information S7.2 and S7.3.
IV.2 Invariant measure for chaotic systems
The concept of an invariant measure originates from the field of ergodic theory and has become a fundamental tool in the study of dynamical systems. First formalized in the context of measure-preserving transformations by Birkhoff and von Neumann in the early 20th century, invariant measures capture the long-term statistical behaviour of systems evolving under deterministic rules [65, 66].
which means that the total probability of the system being in a region remains the same after one step of evolution. This property reflects a kind of statistical balance: although individual trajectories may change, the overall distribution stays stable. In our work, we aim to capture this long-term behaviour by ensuring that the model’s predictions remain consistent with a reference distribution, which we learn using a QCBM.
In turbulent flows, such invariant measures typically manifest as conserved distributions, such as velocity probability, turbulent kinetic energy or energy spectra that encode physical long-term steady states. Capturing these invariant features is essential for ensuring robustness in extrapolative prediction regimes, where a classical machine learning method often suffers from gradient explosion, mode collapse, or long-term error accumulation [67].
IV.3 The quantum circuit Born machine to learn the Q-prior
We develop a quantum-informed learning component in which a QCBM is used to directly learn invariant statistical properties from high-dimensional physical observations. The QCBM defines a parameterised quantum state on an -qubit register, which serves as a generative model over a discretised spatial domain. It does not encode a specific state (e.g., a velocity field) into the circuit. Instead, by observing many snapshots of the system, it learns to reproduce their underlying invariant probability distribution. Specifically, each computational basis state here is mapped bijectively to a spatial coordinate in the simulation grid. The corresponding amplitude is optimised so that its squared magnitude approximates the normalised velocity magnitude at that point. Although the circuit does not encode classical inputs directly via amplitude encoding, it constructs a quantum state whose Born distribution approximates the invariant measure Q-prior of the system. The resulting Born distribution is given by:
where is a parameterised unitary circuit composed of L layers, each consisting of single-qubit rotations and entangling operations. Each computational basis state represents a discrete spatial point, and the quantum amplitudes are trained to reflect the empirical distribution of local velocity magnitudes. This provides a compact, data-driven mapping from physical observables into the Hilbert space, where complex correlations can be more effectively represented. Further details can be found in S1.3 and other sections in the Supplementary Information.
As shown in Algorithm 2, the QCBM is trained as a quantum generative model. Each computational basis state is mapped uniquely to a spatial location in the discretised velocity field, and its target value is the corresponding velocity magnitude . Because the underlying probability distribution over spatial locations is unknown, we adopt a sample–based training strategy that minimises the Maximum Mean Discrepancy (MMD):
Here, denotes the batch of empirical samples drawn from the reference velocity field, while are samples generated by the QCBM. Because the MMD objective is expressed solely through kernel evaluations on finite sample sets, optimisation of the QCBM requires no closed-form expression for the target probability density. This sample-based objective is particularly advantageous for high-dimensional chaotic flows, where the invariant measure is available only as a time-ordered set of simulation snapshots rather than an explicit probability density. By aligning the QCBM’s samples with these empirical trajectories, the model infers the underlying measure without ever needing a closed-form expression.
The trained quantum distribution thus can serve as a data-driven prior capturing the invariant measure of the underlying physical process. Unlike classical regularisers based on fixed heuristics, the QCBM constructs this Q-prior through optimisation on empirical data.
All quantum circuits are implemented using hardware-efficient structures, which are detailed in Supplementary Information S2.2. Training is carried out on noiseless simulators and on superconducting quantum hardware, with circuit transpilation, error mitigation, and device-specific calibration performed using the Qiskit runtime. Further details and the optimisation methods are provided in Supplementary Information S1.3 & S2.3.
IV.4 Quantum hardware implementation
The QCBM model was trained using both emulators and real superconducting quantum processors. To ensure effective execution on near-term quantum hardware and to manage the impact of noise on the quality of Q-priors, this study constrains the total number of trainable parameters and adopts conservative circuit configurations. Specifically, the number of qubits is limited to 10–15, with circuit depths ranging from 3 to 8 layers, resulting in fewer than 300 trainable parameters in total (detail see: Supplementary Information S2.2-2.3). Hardware experiments were mainly conducted on IQM’s 20-qubit superconducting chip (Garnet) [68, 69], using qubits 1–10 as the active subset. Further discussion on hardware sensitivity is provided in Supplementary Sections S2.4 and S2.5. To ensure scalability, the quantum backend was integrated into a GPU-accelerated classical training loop on NVIDIA A100 nodes provided by BEAST GPU cluster (Leibniz Supercomputing Centre). Gradient evaluation, loss computation, and optimiser updates were performed classically and synchronised with the quantum hardware at each iteration. This setup constitutes an applied demonstration of embedding NISQ-era QPUs into scientific machine learning pipelines.
The employed QCBM circuit architecture consisted of alternating layers of parameterised rotations and hardware-native entangling gates. Importantly, the same circuit structure was used across all experiments without dataset-specific tailoring, ensuring general applicability. Although performance could potentially be improved through customised ansatz design, our results demonstrate that the proposed quantum-informed framework is compatible with any hardware-efficient quantum circuit. This adaptability enables the integration of diverse circuit architectures without modifying the overall training pipeline, suggesting broad utility across problem domains and quantum backends.
Each circuit was sampled with 20,000 shots per iteration, and readout errors were mitigated using IQM’s native transpiler and the Mthree measurement mitigation toolkit [70]. Optimisation was performed using L-BFGS [71], Adam, and COBYLA [72], selected based on convergence stability and noise robustness. To prevent mode collapse, such as excessive sampling concentration on the state, runtime monitoring and redundancy checks were incorporated into the training loop. Despite operating with only 10 to 15 qubits, the QCBM successfully captured the structure of invariant distributions over a - to -dimensional Hilbert space, demonstrating the effectiveness of NISQ-compatible priors in modelling high-dimensional systems. Further discussions of the error mitigations, experiment results and ablations are provided in the Supplementary Information S2.5 and S7.
IV.5 Quantum memory efficiency
A defining feature of the proposed framework is its capacity to encode high-dimensional physical information using significantly fewer trainable parameters than conventional deep learning models. The QCBM, operating on only 10–15 qubits and comprising fewer than 300 trainable parameters, successfully captures invariant distributions over spatial domains with dimensionality exceeding . In contrast, a classical neural network capable of learning an equally expressive distribution over the same domain would require at least parameters, assuming a minimal fully connected architecture with comparable granularity. For example, on a grid of spatial points, the lightest fully-connected (FC) network that can approximate an arbitrary probability map with standard universal-approximation accuracy must allocate weights[73, 74]. Even with an aggressively thinned hidden layer this still amounts to tunable parameters—three orders of magnitude more than the rotation angles used by our ten-qubit QCBM. Moreover, classical models lack access to quantum entanglement, limiting their ability to compactly encode non-local correlations and higher-order statistical dependencies prevalent in turbulent or chaotic systems.
For the turbulent–channel dataset, retaining all mid-plane snapshots in double precision requires roughly . Archiving the corresponding QCBM checkpoint for each snapshot ( per file) occupies only , a compression greater than two orders of magnitude while preserving the relevant flow statistics. Comparable savings are obtained for Kolmogorov flow ( parameters, per file, in total) and Kuramoto–Sivashinsky data ( parameters, per file, in total). Full storage calculations are shown and reported in Supplementary Information S8. Such dramatic savings make the Q-prior attractive whenever data are expensive to generate or archive, while still providing a statistically faithful, low-parameter regulariser for the classical Koopman surrogate.
The quantum module’s ability to generate a faithful statistical prior with such a compact parameterisation highlights its potential as a highly efficient generative component within hybrid learning pipelines. This parameter efficiency is particularly relevant in scientific domains where data are sparse or expensive to obtain and store, and where encoding physical constraints is essential for generalisability.
V Data availability
All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Information S4. The training/validation/testing datasets for all three cases are available on https://doi.org/10.5281/zenodo.16419085.
VI Code availability
The complete training code for the QCBM module and the Koopman surrogate model is openly available at the https://github.com/UCL-CCS/QIML.git.

### v3

@@ G.5 Integration of the Traditional Model and the Quantum Prior Guidance
and
with denoting the empirical distribution derived from the predicted field .
While this formulation successfully captures the global distributional shape, it often converges to a low-variance mean-field solution, insufficient to preserve local or high-frequency structures critical for short-term dynamics. To mitigate this, we introduce two additional constraints.
First, we add a MMD term to better match higher-order statistics:
where is a feature mapping into a reproducing kernel Hilbert space.
Second, in some cases (TCF), we also incorporate a peak loss to explicitly align dominant modes of the predicted and target distributions:
where TopK selects the highest-probability bins from both histograms.
The final training loss is a weighted combination of all four components:
This composite objective ensures that predictions remain consistent with the invariant statistics of the underlying physical system, while also preserving critical local features necessary for short- and mid-term forecasting. Our results demonstrate that Q-Priors can significantly enhance the stability and accuracy of machine learning models in high-dimensional chaotic regimes.
Fig. 14 presents the spatial distribution of the normalized velocity error across the domain under three modeling regimes. The top panel displays error patterns obtained without Q-Priors, revealing widespread, spatially incoherent deviations with peak errors exceeding 10%. Incorporating a Q-Prior from a classical emulator (middle panel) leads to a pronounced reduction in error magnitude and spatial variance, with the majority of regions falling below 4%. Further improvement is observed when employing the Q-Prior trained on real quantum hardware (bottom panel), where error distributions are both quantitatively reduced and exhibit greater spatial smoothness, indicating enhanced generalization and stability in chaotic regime reconstruction. Fig. 15 shows the TCF velocity magnitude distribution predicted under the same three learning regimes (no Q-Prior, emulator-based Q-Prior, and hardware-based Q-Prior), compared to the empirical target distribution. Without quantum guidance, the model fails to capture the dominant statistical mode—both the modal peak and its amplitude are significantly misaligned. With simulator-based Q-Prior, the predicted distribution nearly perfectly matches the raw data in both location and shape, confirming the Q-Prior’s capacity to encode high-dimensional invariant measures with minimal parameterization. Hardware-based Q-Prior introduces moderate deviations, including a downward shift in peak probability and a small modal displacement; however, the distribution remains substantially closer to the ground truth than the classical result.
System | Qubits | Trainable | Raw data | Q-Prior file | Compression | Device |
parameters | (full set) | (full set) | (ratio) | |||
Kuramoto–Sivashinsky | 10 | Emulator | ||||
2D Kolmogorov flow | 10 | Emulator | ||||
Turbulent channel flow | 15 | Emulator | ||||
Turbulent channel flow | 10 | IQM-Garnet |
Appendix H Quantum Parameter Efficiency and Quantum Memory Advantage
This section explain the parameter efficiency and memory advantage for the QIML. According to the parameter numbers in this work, as shown in table 4, we used 120 parameters on the KS system, 180 on the Kolmogorov flows on the emulator, and 240 or 300 for the TCF system on the hardware and emulator. As is evident, our method uses a remarkably small number of parameters to learn the necessary statistical information. We have already shown the orders-of-magnitude difference in parameter counts compared to classical models in table 3. Both the parameter efficiency discussed here and the memory advantage that follows originate from the same foundational quantum advantage [65]: the ability of a quantum circuit to logarithmically compress classical data. This is the fundamental source of the quantum advantage established in our work, a topic we have discussed in detail in Sec. VI. The critical challenge in harnessing this advantage is circumventing the limitations imposed by Holevo’s bound. This bound arises because retrieving the complete classical information from the quantum state would necessitate an exponential number of measurements, rendering the approach impractical. This obstacle is overcome because the task of guiding a dynamical model does not necessitate access to the complete classical dataset. Instead, the framework’s objective is to extract a statistical summary, the distribution of velocity information, which represents the system’s invariant measure. For a chaotic system, such a measure inherently contains far less information than the full dataset from which it is derived. The ability to function with only this compressed distributional information allows the method to bypass the need for an exponential number of measurements. This principle is the foundation for the quantum memory advantage, which is now detailed for the three systems investigated in this paper.
H.1 Kuramoto–Sivashinsky Equation
The raw KS data comprise trajectories, each recorded for time steps on a -point spatial grid, totalling GB. For training we down-sample to trajectories, each with temporal frames and spatial points, leading to a working tensor of shape per trajectory. With double precision ( bytes per value) a single trajectory occupies
Thus, the full down-sampled dataset requires . Each trajectory is paired with an individual quantum generator containing qubits, circuit layers and rotation parameters, for a raw weight file of bytes; including metadata the stored checkpoint is kB. Archiving one checkpoint per trajectory, therefore, costs . Consequently, the learned Q-Priors reduce storage from to , a compression factor of around while still capturing the essential statistics used by the classical machine learning model.
H.2 2D Kolmogorov Flows
The total dataset comprises trajectories, each containing temporal snapshots on a grid. For a single velocity component, one snapshot occupies
so that one complete trajectory requires and the full set of trajectories stores . The corresponding quantum generator employs qubits, layers and trainable rotation angles. Saved in double precision, one checkpoint is . Archiving a separate checkpoint for every snapshot therefore costs
yielding a compression factor of
while faithfully reproducing the empirical invariant measure.
H.3 Turbulent Channel Flow
For the TCF benchmark, each snapshot is stored as a array of double-precision values (8 bytes per entry). A single velocity component therefore, occupies
so that the full sequence of 595 snapshots amounts to
Retaining all three velocity components raises the per-trajectory footprint to roughly 500 MB.
The quantum generators used to reproduce the same invariant statistics each employ qubits, layers, and 240 rotation angles. Each saved quantum generator checkpoint, with minimal metadata, is about 4.3 kB. Given that 595 such quantum generator models are retained (one for each snapshot), the total storage for the entire set of quantum generator checkpoints amounts to . Relative to the raw three-component trajectory (), this represents a compression factor of approximately 200, which is over two orders of magnitude, while still capturing the statistics required by the classical Koopman machine learning model.
Appendix I Performance and Statistical Metrics
This section provides the mathematical definitions for the primary metrics used in this paper to evaluate model performance and analyse the system’s statistical properties.
I.1 Temporal Autocorrelation
The temporal autocorrelation is used to measure the memory of the dynamical system, quantifying the correlation of a time series with a delayed version of itself. For a discrete time series (the value at a specific spatial point at time step ), with mean and total length , the normalized temporal autocorrelation at a time lag of steps is defined as:
I.2 Error Metrics
To quantify the difference between the ground truth data and the model predictions, the following error metrics are used.
The absolute error provides a direct, point-wise measure of the deviation between a predicted field and the ground truth field :
To specifically quantify how well the model reproduces the temporal correlation structure, we define the normalized relative auto-correlation error . This measures the relative L2 error between the true autocorrelation function and the predicted one :